You will receive 100 British bonds that pay interest forever. The amount of annual interest payments that you we ece ve is S5.000. If you could invest yOur money at 4.25%, how much are these bonds worth today? O A $55,000 B. $197,250 C. $250,000 O D. $117.647 ? E. $64,480 are considering the purchase of XYZ Companys common stock which will pay a $1.00 per share dividend one year from the date of purchase. The divid end is expected to grow at the rate of 4% per year if the appropriate discount rate for this investment is 14%, what is the price of one share of this stock A. $7.14 ?. $25.00 O c. $10.00 O D. Cannot be determined without maturity date retirement plan guarantees to pay you or your estate a fixed amount for 20 years. Al the time of retirement you will benefits. How much will your annual benafits be., assuming the first have $31.360 to your credit in the plan. Theplan ant pates earning 8% interest annually over the pero d you receive payment occurs one year from your retirement date? O A. $2,000 O B. $6,272 O C. $3,194 O D. $682 19 29. What is the present value of an annuity of $12 received at the end of each year for seven years? Assume a discount rate of 11%. The first payment will be received one year from today (round to the nearest $1). O A $118 ??. $40 O c. $25 OD. $57 20. An investment has a nominal interest rate of 12% annually, but interest on the investment is compounded monthly Therefore, the annual compounding rate of returnis: 30. OA. 13.89%. O B. 1296. ??. 12.68%. O D. 12.36%
Answer #1

26.

bonds worth today is

=5000/4.25%

=117647

27.

price of one share=1/(14%-4%)=10.00

28.

annual benefits

=31360/((1-(1+8%)^(-20))/8%)

=3194

29.

present value

=12*((1-(1+11%)^(-7))/11%)

=57

30.

annual compounding returns is=(1+(12%/12))^12-1

=12.68%

the above is answer..